# Marrying Autoregressive Transformer and Diffusion with Multi-Reference Autoregression

### v1

@@ Abstract
We introduce TransDiff, the first image generation model that marries Autoregressive (AR) Transformer with diffusion models. In this joint modeling framework, TransDiff encodes labels and images into high-level semantic features and employs a diffusion model to estimate the distribution of image samples. On the ImageNet $256\times 256$ benchmark, TransDiff significantly outperforms other image generation models based on standalone AR Transformer or diffusion models. Specifically, TransDiff achieves a Fréchet Inception Distance (FID) of 1.61 and an Inception Score (IS) of 293.4, and further provides ×$\times 2$ faster inference latency compared to state-of-the-art methods based on AR Transformer and ×$\times 112$ faster inference compared to diffusion-only models. Furthermore, building on the TransDiff model, we introduce a novel image generation paradigm called Multi-Reference Autoregression (MRAR), which performs autoregressive generation by predicting the next image. MRAR enables the model to reference multiple previously generated images, thereby facilitating the learning of more diverse representations and improving the quality of generated images in subsequent iterations. By applying MRAR, the performance of TransDiff is improved, with the FID reduced from 1.61 to 1.42. We expect TransDiff to open up a new frontier in the field of image generation.
Figure 1: Generated samples from TransDiff trained on ImageNet. Top: $512\times 512$ and $256\times 256$ samples. Middle: effect of semantic feature diversity in TransDiff on image quality(left to right: increasing diversity.). Bottom: results of semantic features fusion from images of different classes. (The first two columns show images from two classes; the third shows the fused result.)
@@ 1 Introduction
In recent years, image generation models have made significant progress driven by transformers vaswani2017attention ; esser2021taming ; van2017neural ; razavi2019generating and diffusion models ho2020denoising ; dhariwal2021diffusion ; nichol2021improved ; peebles2023scalable ; rombach2022high ; chen2024pixart ; flux2024 . Currently, the two main approaches are AR Transformer and diffusion models. AR Transformer use tokenized image representations that convert images from pixel space into discrete visual tokens through vector quantization (VQ) van2017neural for AR image generation. In contrast, diffusion models employ a progressive denoising process, starting from random Gaussian noise and ultimately generating images.
Although these two approaches have achieved remarkable success in image generation, they face fundamental limitations inherent in their design. On the one hand, a key challenge for AR Transformer image generation is the $O(n^{2})$ computational complexity of causal attention, which becomes especially prohibitive for high-resolution images. To mitigate this, most current AR-based methods convert raw pixels into a latent space representation—e.g., $32\times 32$ instead of the original $256\times 256\times 3$ image space. This compressed representation enables efficient model training and inference. However, the high-rate compression approach, while mitigating the complexity of $O(n^{2})$ through discrete representations esser2021taming ; van2017neural ; razavi2019generating learned in the latent space, can compromise image quality by reducing detail fidelity. The method essentially achieves computational efficiency at the cost of information loss, as it quantizes continuous features and trains on discrete tokens. On the other hand, diffusion models can represent arbitrary probability distributions, enabling the generation of high-quality, photorealistic images. However, their iterative denoising process inherently results in slower inference speeds compared to AR approaches. Moreover, diffusion models only utilize a frozen LM raffel2020exploring to represent text condtition or employ an n-layer MLP to encode class condtition, which limits their representational capacity.
@@
A natural question arises: Can we marry the strengths of AR Transformers and diffusion models into a unified architecture? We propose TransDiff, a novel approach that integrates the advantages of both AR Transformers and diffusion models within a unified training framework. Specifically, TransDiff first employs an AR Transformer to encode inputs into high-level semantic features, then uses a diffusion model (i.e., the "diffusion-based decoder") to decode these features into images. Through joint training, the AR Transformer learns to extract high-level semantics while the diffusion-based decoder develops the capability to interpret these semantic features precisely (e.g., Fig. 1).
@@
Current AR models can be broadly categorized into two types: token-level AR models ramesh2021zero ; cao2023efficient ; chen2020generative ; van2017neural ; van2016conditional and scale-level AR models tian2024visual ; liu2024alleviating ; ma2024star ; zhang2024var . Token-level models generate images by predicting one token at each step, with the model only observing an incomplete set of image tokens during generation. In contrast, scale-level models generate images by progressively predicting the next scale, where the model refers to a coarse and blurry image version at each step. In both cases, the predictions rely on incomplete or imprecise information, which limits the overall quality and diversity of the generated images.
@@
To address these limitations, we propose a novel autoregressive generation paradigm—Multi-Reference AutoRegression (MRAR)—where the model predicts a complete image at each step. This approach enables the model to leverage more comprehensive and complete reference information during generation, resulting in more diverse and higher-quality image outputs (e.g., Fig. 2).
@@
We validate the effectiveness of TransDiff with MRAR as its autoregressive strategy on the challenging ImageNet deng2009imagenet benchmark. This approach achieves a 1.42 Fréchet Inception Distance (FID) score, outperforming diffusion models with the equal parameter count (FID $1.58$). While maintaining inference speed comparable to state-of-the-art AR Transformer models, it significantly outperforms diffusion models in inference efficiency.
Figure 2: Compare of Token-Level AR, Scale-Level AR and MRAR.
@@ 2.1 Diffusion Models for Image Generation
Diffusion models gradually remove noise from images to restore them to clear images. Early works used U-Net as the denoiser HoSCFNS22 ; RombachBLEO22 ; ho2020denoising ; DBLP:conf/iclr/LiuG023 ; DBLP:conf/iclr/0011SKKEP21 . Such as, CDM HoSCFNS22 uses cascaded U-net diffusion models to generate high-fidelity images. And, LDM RombachBLEO22 performs the diffusion processes in the latent space of pretrained autoencoders and significantly reduces the computational cost. Recently, with the significant progress of the transformer architecture in the field of NLP, a growing number of works have begun using transformers as denoisers for image generation PeeblesX23 ; DBLP:conf/icml/EsserKBEMSLLSBP24 ; DBLP:journals/corr/abs-2405-08748 ; flux2024 , such as, DiT PeeblesX23 is the first work to use the transformer architecture instead of the previous U-Net architecture and to explore the scalability of diffusion transformers. And, Esser et al. DBLP:conf/icml/EsserKBEMSLLSBP24 improve rectified flow models with perceptually-biased noise sampling and introduces a novel dual-modality transformer.
@@ 2.2 Autogressive Models for Image Generation
Autoregressive models generate images by regressively predicting the next token LeeKKCH22 ; ChangZJLF22 ; DBLP:journals/corr/abs-2502-20388 ; DBLP:journals/corr/abs-2411-00776 ; esser2021taming or next scale tian2024visual ; liu2024alleviating ; ma2024star ; zhang2024var . For next token prediction, RQ-Transformer LeeKKCH22 utilizes a Residual-Quantized VAE (RQ-VAE) for autoregressive image generation, resulting in lower computational costs and faster sampling speeds. Maskgit ChangZJLF22 utilizes a bidirectional transformer decoder that learns to predict randomly masked image tokens by attending to tokens in all directions, significantly enhancing the efficiency of autoregressive decoding. For next scale prediction, VAR tian2024visual proposes a coarse-to-fine "next-scale prediction" approach to redefine autoregressive learning for images, enabling GPT-style AR models to surpass diffusion transformers in image generation for the first time. DiMR liu2024alleviating incrementally enhances features at various scales, enabling a gradual improvement in visual detail from coarse to fine levels.
@@
3 Approach
@@ 3.1 TransDiff
Figure 3: Our Transdiff approach overview. Wherein (a) and (c) are the Vanilla Autoregressive Transformer and the Vanilla Diffusion Model, respectively. (b) is our TransDiff approach. (d) and (e) provide a detailed explanation of 1-step AR and MRAR.
We introduce TransDiff, which first unifies two dominant methods in contemporary image generation: the AR Transformer and the diffusion model. In brief, we employ the AR Transformer to encode input into high-level semantic features serving as a conditioning signal, and utilize a diffusion decoder peebles2023scalable to generate the image. This section begins with a brief overview of these two methodologies, followed by a detailed explanation of how we systematically integrate them into a unified framework.
AR Transformer. Conventional autoregressive image generation methods typically rely on a pre-trained tokenizer to quantize images into discrete tokens, which are then used to reconstruct the original image. However, the limited capacity of the VQ model imposes a fundamental constraint on the reconstruction quality . This bottleneck arises from the inherent trade-off between compression ratio and information retention: a smaller compression ratio increases computational complexity, while a larger compression ratio leads to significant loss of information during quantization.
Unlike common AR Transformer models radford2019language ; brown2020language , which take discrete tokens as input, we let $x=(x_{1},x_{2},\ldots,x_{N})$ denote a sequence of continuous features, where $x_{n}\in\mathbb{R}^{l\times d}$ The AR Transformer models assume that the probability of each $x_{n}$ depends exclusively on its preceding tokens $(x_{1},x_{2},\ldots,x_{n-1})$. This assumption of unidirectional dependency enables the joint probability of the entire sequence to be factorized as a product of conditional probabilities:
@@
The loss function of AR Transformer models is typically defined as:
$\mathcal{L}_{AR}={min}\frac{1}{N}\sum_{n=0}^{N}\mathrm{LF}(x_{n},\text{AR Transformer}(x_{0},x_{1},...,x_{n-1}))$ (2)
@@
Diffusion Decoder. We adopt DiT peebles2023scalable as our diffusion decoder. In addition, we replace the original DDPM ho2020denoising objective with Flow Matching DBLP:conf/iclr/LipmanCBNL23 . Flow Matching learns a transformation from latent variables $\epsilon\sim\mathcal{N}(0,1)$ to samples $z$ drawn from the data distribution $q$, formulated through an ordinary differential equation (ODE):
@@
$\mathcal{L}_{diff}=\min\mathbb{E}_{t\sim Uniform([0,1]),x\sim P,c\sim\text{ARTransformer}}\left[||(\epsilon-x)-\psi_{\theta}(x^{t},\ t,\ c)||^{2}\right]$ (5)
@@
Marrying AR Transformer with Diffusion Decoder. In this paper, we introduce TransDiff, which, to the best of our knowledge, is the first framework that enable joint training of the AR Transformer and the diffusion model(i.e., diffusion decoder). While AR Transformer-based image generation methods use VQ-VAE to map images into discrete tokens, our approach employs a VAE to project images into a continuous latent space. Subsequently, our AR Transformer encodes the obtained image latent into high-level semantic features.
$c=\text{AR Transformer}(\text{input})$ (6)
where $\text{input}\in\mathbb{R}^{(h\times w)\times d}$ denotes the the continuous latent of image, $c\in\mathbb{R}^{(h/f\times w/f)\times(d\times f\times f)}$ denotes the condition for diffusion decoder, where $f$ denotes the compression ratio.
Then, these features serve as conditioning inputs to the diffusion decoder to generate the image. Detailed as follows:
output =$\text{output }=\text{Diff}(c,\epsilon)$ (7)
where Diff denotes the Diffusion Deocder, $\epsilon\sim\mathcal{N}(0,I)$ and $output\in R^{(h\times w)\times d}$ denotes the the continuous latent of output image.
Finally, by combining the loss of AR Transformer model (i.e.,Eq. 2 )and the diffusion decoder(i.e., Eq. 5), the overall loss of our approach can be defined as:
$\mathcal{L}_{\text{all}}=\min\frac{1}{N}\sum_{n=0}^{N}\mathbb{E}_{\begin{subarray}{c}t\sim\text{Uniform}([0,1]),x\sim P\end{subarray}}\Bigg{[}$ $\Big{\|}(\epsilon-x_{n})-$ (8)
$\psi_{\theta}(x_{n}^{t},\ t,\ \text{AR Transformer}(x_{0},x_{1},\dots,x_{n-1}))\Big{\|}^{2}\Bigg{]}$
where $x_{n}$ denotes the $n$-th known input latent for AR Transformer, and $x_{n}^{t}$ represents the diffusion sample of $x_{n}$ at time step $t$.
@@ 3.2 Multi-Reference Autoregressive (MRAR)
In the autoregressive component of our approach, we introduce a novel autoregressive generation paradigm, Multi-Reference Autoregressive (MRAR), which diverges from traditional autoregressive generation paradigm. Unlike the standard one-step autoregressive (1-step AR) method, which generates predictions sequentially from the previous output, MRAR incorporates multiple references from the entire sequence of previous outputs. Our experimental results demonstrate that MRAR consistently outperforms the one-step autoregressive strategy, offering significant improvements in performance. Based on these observations, we adopt MRAR as the preferred autoregressive strategy in our model, positioning it as a new paradigm for autoregressive modeling.
@@
As shown in the Figure 3, the goal of 1-Step AR encode high-level semantic features $c$ through a single forward pass of the AR transformer, conditioned on the class label, which is then decoded by the Diffusion Decoder to generate the output.
Specifically, the input can be formalized as follows:
$\text{Mask}=\text{Concat}([\text{m},\text{m},...,\text{m}])$ (9)
@@
where, C denotes the class embedding of image, and m represents the mask embedding. And $\text{C}\in\mathbb{R}^{64\times d}$, $\text{Mask}\in\mathbb{R}^{(h\times w)\times d}$.
Then, we follow the forward propagation and loss computation of TransDiff:
[…]
@@
Furthermore, as shown in the Figure 3 (d), in 1-Step AR, the attention mask of the AR Transformer in 1-Step AR is a bidirectional attention matrix filled with zeros.
The inference process of 1-Step AR is shown in the Algorithm 1.
3.2.2 Multi-Reference Autoregressive (MRAR)
In the 1-Step AR setting, we observe that in the image features $c$ (as defined in the Eq. 10) of a category, the more diverse the features relevant to that category, the better the quality of the generated images (A detailed experimental analysis is provided in the Sec. 4.2.2). Based on the above observations, we propose MRAR, which learns more comprehensive features by referencing multiple ’past’ generated images. This allows it to generate higher-quality images through iterations(A detailed experimental analysis is provided in Sec. ). The detailed process is shown in the Figure 3 (e). Therefore, building on the training of 1-Step AR, we further fine-tune the model using MRAR.
Specifically, the input can be formalized as follows:
$\text{input}=\text{Concat}([\text{C},\text{Mask},x_{\text{img}_{0}},x_{\text{img}_{1}},...,x_{\text{img}_{n-1}}])$ (11)
where $\{x_{\text{img}_{i}}\}_{i=1}^{n}\in\mathbb{R}^{(h/f\times w/f)\times(d\times f\times f)}$ denotes the latent of the $i$-th image under the same label, and $f$ is the compression ratio. $\text{Mask}\in\mathbb{R}^{(h/f\times w/f)\times(d\times f\times f)}$ represents the mask embedding.
Then, we follow the forward propagation and loss computation of TransDiff:
$c=\text{AR Transformer(input)}=\text{Concat}([c_{\text{img}_{0}},c_{\text{img}_{1}},c_{\text{img}_{2}},...,c_{\text{img}_{n}}])$ (12)
$\text{output}=\text{Diff}(c,\epsilon)=\text{Concat}([o_{\text{img}_{0}},o_{\text{img}_{1}},o_{\text{img}_{2}},...,o_{\text{img}_{n}}])$
where $\{c_{\text{img}_{i}}\}_{i=0}^{n}\in\mathbb{R}^{(h/f\times w/f)\times(d\times f\times f)}$ and $\{o_{\text{img}_{i}}\}_{i=0}^{n}\in\mathbb{R}^{(h\times w)\times d}$ denote the condition features and the final output of the $i$-th image, respectively. Moreover, the final generated image is denoted as $o_{\text{img}_{n}}\in\mathbb{R}^{(h\times w)\times d}$. As illustrated in the Figure 3 (e), the attention mask in the AR Transformer of MRAR is implemented as a causal attention matrix.
The inference process of MRAR is shown in the Algorithm 2.

### v3

@@ Abstract
We present TransDiff, the first image generation model that marries Autoregressive (AR) Transformer with diffusion models. In this unified framework, TransDiff encodes labels and images into high-level semantic features and employs a diffusion model to estimate the distribution of image samples. On the ImageNet $256\times 256$ benchmark, TransDiff significantly outperforms all other image generation models based on standalone AR Transformer or diffusion models. Specifically, TransDiff achieves a Fréchet Inception Distance (FID) of 1.61 and an Inception Score (IS) of 293.4, and further provides $\times 2$ faster inference latency compared to state-of-the-art methods based on AR Transformer and $\times 112$ faster inference compared to diffusion-only models. Moreover, building on the TransDiff model, we introduce a novel image generation paradigm, namely Multi-Reference Autoregression (MRAR), which performs autoregressive generation by predicting the next image. It enables the model to reference multiple previously generated images, thereby facilitating the learning of more diverse representations and improving the quality of generated images in subsequent iterations. By applying MRAR, the performance of TransDiff is improved, with the FID reduced from 1.61 to 1.42. We expect TransDiff to open up a new frontier in the field of image generation. Refer to caption Figure 2: Comparison of Token-Level AR, Scale-Level AR and MRAR paradigms.
Figure 1: Generated samples from TransDiff trained on ImageNet. Top: $512\times 512$ and $256\times 256$ samples. Middle: effect of semantic feature diversity in TransDiff on image quality(left to right: increasing diversity.). Bottom: results of semantic features fusion. (The first two columns show images from two classes; the third shows the fused result.)
@@ 1 Introduction
Image generation has achieved remarkable progress driven by transformer architectures esser2021taming ; van2017neural ; razavi2019generating and diffusion models ho2020denoising ; dhariwal2021diffusion ; nichol2021improved ; peebles2023scalable ; rombach2022high ; chen2024pixart ; flux2024 . Currently, two dominant approaches prevail: Autoregressive (AR) Transformers and diffusion models. AR Transformers employ vector quantization (VQ) van2017neural to convert images into discrete tokens for sequential generation. Diffusion models progressively denoise from random Gaussian noise to generate high-fidelity images.
However, both paradigms face fundamental limitations. AR Transformers suffer from the $O(n^{2})$ computational complexity of causal attention. This complexity becomes prohibitive for high-resolution images. To mitigate this, most AR-based methods tian2024visual ; sun2024autoregressive compress images into discrete latent representations (e.g., $32\times 32$ instead of $256\times 256\times 3$). This compression achieves computational efficiency at the cost of information loss through quantization. The discrete representation bottleneck fundamentally constrains generation quality. Conversely, diffusion models excel at producing high-quality images by modeling arbitrary probability distributions li2024autoregressive . However, their iterative denoising process results in significantly slower inference speeds compared to AR approaches.
@@
A natural question arises: Can we take advantages of both models and have a new paradigm for image generation? To this end, we present TransDiff, the first unified framework that combines AR Transformer with diffusion models. Different from the recent attempt BLIP3-ochen2025blip3 , which relies on two-stage training and sequentially learns AR transformer and diffusion model, our TransDiff is jointly trained with an end-to-end manner. Specifically, TransDiff employs an AR Transformer to encode inputs into high-level semantic features and apply a diffusion model as a decoder to generate images from these semantic representations. Through training, the AR transformer learns to extract precise semantic features. The diffusion decoder develops the capability to interpret these features effectively (as shown in Figure 1 Middle and Bottom).
@@
Building upon our TransDiff, we further investigate how to improve the AR paradigm and achieve better performance. Current autoregressive paradigms can be categorized into two types: token-level AR models ramesh2021zero ; cao2023efficient ; chen2020generative ; van2017neural ; van2016conditional and scale-level AR models tian2024visual ; liu2024alleviating ; ma2024star ; zhang2024var . Token-level models generate images by predicting one token at each step. During generation, they observe only incomplete image tokens. Token-level models observe only incomplete image tokens during generation and scale-level models refer to coarse and blurry image versions at each step. Both paradigms rely on incomplete or imprecise information. This limitation constrains generation quality and diversity.
@@
To bridge this gap, we furthre introduce Multi-Reference Autoregression (MRAR). This is a novel AR generation paradigm where the model predicts a complete image at each step. Then MRAR enables the model to leverage comprehensive reference information from the previous steps during generation, achieving more diverse and higher-quality outputs (e.g., Figure 2).
@@
We validate TransDiff with MRAR on the challenging ImageNet deng2009imagenet benchmark. Our TransDiff achieves a 1.42 Fréchet Inception Distance (FID) score. This outperforms diffusion models with equal parameter count (FID 1.58). Meanwhile TransDiff also maintains inference speed comparable to state-of-the-art AR Transformer models. It significantly outperforms diffusion models in inference efficiency.
@@ 2.1 Diffusion Models for Image Generation
Diffusion models gradually remove noise from images to restore them to clear images. Early works used U-Net as the denoiser HoSCFNS22 ; RombachBLEO22 ; ho2020denoising ; DBLP:conf/iclr/LiuG023 ; DBLP:conf/iclr/0011SKKEP21 . Such as, CDM HoSCFNS22 uses cascaded U-net diffusion models to generate high-fidelity images. And, LDM RombachBLEO22 performs the diffusion processes in the latent space of pretrained autoencoders and significantly reduces the computational cost. Recently, with the significant progress of the transformer architecture in the field of NLP, a growing number of works have begun using transformers as denoisers for image generation PeeblesX23 ; DBLP:conf/icml/EsserKBEMSLLSBP24 ; DBLP:journals/corr/abs-2405-08748 ; flux2024 , such as, DiT PeeblesX23 is the first work to use the transformer architecture instead of the previous U-Net architecture and to explore the scalability of diffusion transformers. And, Esser et al. DBLP:conf/icml/EsserKBEMSLLSBP24 improved rectified flow models with perceptually-biased noise sampling and introduces a novel dual-modality transformer.
@@ 2.2 Autogressive Models for Image Generation
AR models generate images by regressively predicting the next token LeeKKCH22 ; ChangZJLF22 ; DBLP:journals/corr/abs-2502-20388 ; DBLP:journals/corr/abs-2411-00776 ; esser2021taming or next scale tian2024visual ; liu2024alleviating ; ma2024star ; zhang2024var . For next token prediction, RQ-Transformer LeeKKCH22 utilizes a Residual-Quantized VAE (RQ-VAE) for autoregressive image generation, resulting in lower computational costs and faster sampling speeds. Maskgit ChangZJLF22 utilizes a bidirectional transformer decoder that learns to predict randomly masked image tokens by attending to tokens in all directions, significantly enhancing the efficiency of autoregressive decoding. For next scale prediction, VAR tian2024visual proposes a coarse-to-fine "next-scale prediction" approach to redefine autoregressive learning for images, enabling GPT-style AR models to surpass diffusion transformers in image generation for the first time. DiMR liu2024alleviating incrementally enhances features at various scales, enabling a gradual improvement in visual detail from coarse to fine levels.
More recently, BLIP3-ochen2025blip3 relies on two-stage training process: it first uses EVA-CLIP to encode images into continuous visual embeddings, which are then reconstructed using a diffusion model. Then the AR Transformer in BLIP3-o is trained to generate these visual embeddings. However, the EVA-CLIP’s original training objective is not aligned with generation tasks, thus the visual embeddings are not optimal for the diffusion models.
@@
Figure 3: Overview of our Transdiff. Wherein (a) and (c) are the Vanilla Autoregressive Transformer and the Vanilla Diffusion Model, respectively. (b) is our TransDiff approach. (d) and (e) provide a detailed explanation of 1-step AR and MRAR.
3 Methodology
@@ 3.1 TransDiff
We introduce TransDiff, which first unifies two dominant methods in contemporary image generation: the AR Transformer and the diffusion model. In brief, we employ the AR Transformer to encode input into high-level semantic features serving as a conditioning signal, and utilize a diffusion decoder peebles2023scalable to generate the image. This section begins with a brief introduction of these two paradigms, followed by a detailed explanation of how we systematically integrate them into a unified framework.
3.2 AR Transformer (ART)
Vanilla autoregressive image generation methods typically rely on a pre-trained tokenizer to quantize images into discrete tokens, which are then used to reconstruct the original image. However, the limited capacity of the VQ model imposes a fundamental constraint on the reconstruction quality . This bottleneck arises from the inherent trade-off between compression ratio and information retention: a smaller compression ratio increases computational complexity, while a larger compression ratio leads to significant loss of information during quantization.
Unlike common AR Transformer models radford2019language ; brown2020language , which take discrete tokens as input, we let $x=(x_{1},x_{2},\ldots,x_{N})$ denote a sequence of continuous features, where $x_{n}\in\mathbb{R}^{l\times d}$ . The AR Transformer models assume that the probability of each $x_{n}$ depends exclusively on its preceding tokens $(x_{1},x_{2},\ldots,x_{n-1})$. This assumption of unidirectional dependency enables the joint probability of the entire sequence to be factorized as a product of conditional probabilities:
@@ 3.2 AR Transformer (ART)
The loss function is typically defined as:
$\mathcal{L}_{\textit{AR}}=\frac{1}{N}\sum_{n=0}^{N}\mathrm{LF}(x_{n},\text{ART}(x_{0},x_{1},...,x_{n-1}))$ (2)
@@
3.3 Diffusion Decoder
We adopt DiT peebles2023scalable as our diffusion decoder. In addition, we replace the original DDPM ho2020denoising objective with Flow Matching DBLP:conf/iclr/LipmanCBNL23 . Flow Matching learns a transformation from latent variables $\epsilon\sim\mathcal{N}(0,1)$ to samples drawn from the data distribution, formulated through an ordinary differential equation (ODE):
@@ 3.3 Diffusion Decoder
$\begin{split}\mathcal{L}_{\text{diff}}=\ &\mathbb{E}_{\begin{subarray}{c}t\sim\text{Uniform}([0,1])\\ x\sim P\\ c\sim\text{ART}\end{subarray}}\left[\left\|(\epsilon-x)-\psi_{\theta}\left(x^{t},\ t,\ c\right)\right\|^{2}\right]\end{split}$ (5)
@@
3.4 The Final TransDiff Architecture
We present TransDiff, which serves as the first framework that enable joint training of the AR Transformer and the diffusion model(i.e., diffusion decoder). While AR Transformer-based image generation methods use VQ-VAE to map images into discrete tokens, our approach employs a VAE to project images into a continuous latent space. Subsequently, our AR Transformer encodes the obtained image latent into high-level semantic features.
$c=\text{ART}(\text{input})$ (6)
where $\text{input}\in\mathbb{R}^{(h\times w)\times d}$ denotes the the continuous latent of image, $c\in\mathbb{R}^{(h/f\times w/f)\times(d\times f\times f)}$ denotes the condition for diffusion decoder, where $f$ denotes the compression ratio. Then, these features serve as conditioning inputs to the diffusion decoder to generate the image. Detailed as follows:
$\text{output }=\text{Diff}(c,\epsilon)$ (7)
where Diff denotes the Diffusion Deocder, $\epsilon\sim\mathcal{N}(0,I)$ and $output\in\mathbb{R}^{(h\times w)\times d}$ denotes the the continuous latent of output image. Finally, by combining the loss of AR Transformer model ( i.e., Eq. 2 )and the diffusion decoder( i.e., Eq. 5), the overall loss function can be defined as:
$\mathcal{L}_{\text{all}}=\frac{1}{N}$ $\sum_{n=0}^{N}\mathbb{E}_{\begin{subarray}{c}t\sim\text{Uniform}([0,1]),x\sim P\end{subarray}}\Bigg{[}\Big{\|}(\epsilon-x_{n})$- (8)
$\psi_{\theta}(x_{n}^{t},\ t,\ \text{ART}(x_{0},x_{1},\dots,x_{n-1}))\Big{\|}^{2}\Bigg{]}$
where $x_{n}$ is the $n$-th input latent for AR transformer, and $x_{n}^{t}$ represents the diffusion sample of $x_{n}$ at time step $t$.
@@ 3.5 Multi-Reference Autoregressive (MRAR)
In the AR component of TransDiff, we introduce a novel AR generation paradigm, Multi-Reference Autoregressive (MRAR), which diverges from traditional AR generation paradigm. Unlike the standard 1-step AR method, which generates predictions sequentially from the previous output, MRAR incorporates multiple references from the entire sequence of previous outputs. Our experimental results in Table 4 demonstrate that MRAR consistently outperforms the one-step AR strategy, offering significant improvements in performance. Based on these observations, we adopt MRAR as the preferred AR strategy in our model, positioning it as a new paradigm for AR modeling.
@@
As in the Figure 3(d), the goal of 1-Step AR encode the input high-level into semantic features $c$ through a single forward pass of the AR transformer, conditioned on the class label, which is then decoded by the diffusion decoder to generate the output. Specifically, the input is:
$\text{Mask}=\text{Concat}([\text{m},\text{m},.,\text{m}])$ (9)
@@
where, C denotes the class embedding of image, and m represents the mask embedding. And $\text{C}\in\mathbb{R}^{64\times d}$, $\text{Mask}\in\mathbb{R}^{(h\times w)\times d}$. Then, we follow the forward propagation and loss computation of TransDiff:
$c=\text{ART(input)}$ (10)
@@
Furthermore, as shown in the Figure 3 (d), in 1-Step AR, the attention mask of the AR Transformer in 1-Step AR is a bidirectional attention matrix filled with zeros. The inference process of 1-Step AR is shown in the Appendix.
3.5.2 The MRAR Paradigm
In the 1-Step AR setting, we observe that in the image features $c$ (as defined in the Eq. 10) of a category, the more diverse the features relevant to that category, the better the quality of the generated images (A detailed experimental analysis is provided in the Figure. 5). Based on the above observations, we propose MRAR, which learns more comprehensive features by referencing multiple generated images in the previous steps. This allows it to generate higher-quality images through iterations . The detailed process is shown in the Figure 3 (e). Therefore, building on the training of 1-Step AR, we further fine-tune the model using MRAR. Specifically, the input can be formalized as follows:
$\text{input}=\text{Concat}([\text{C},\text{Mask},x_{\text{img}_{0}},x_{\text{img}_{1}},.,x_{\text{img}_{n-1}}])$ (11)
where $\{x_{\text{img}_{i}}\}_{i=1}^{n}\in\mathbb{R}^{(h/f\times w/f)\times(d\times f\times f)}$ denotes the latent of the $i$-th image under the same label, and $f$ is the compression ratio. $\text{Mask}\in\mathbb{R}^{(h/f\times w/f)\times(d\times f\times f)}$ represents the mask embedding. Then, we follow the forward propagation and loss computation of TransDiff:
$c=\text{ART(input)}=\text{Concat}([c_{\text{img}_{0}},c_{\text{img}_{1}},c_{\text{img}_{2}},.,c_{\text{img}_{n}}])$ (12)
$\text{output}=\text{Diff}(c,\epsilon)=\text{Concat}([o_{\text{img}_{0}},o_{\text{img}_{1}},o_{\text{img}_{2}},.,o_{\text{img}_{n}}])$
where $\{c_{\text{img}_{i}}\}_{i=0}^{n}\in\mathbb{R}^{(h/f\times w/f)\times(d\times f\times f)}$ and $\{o_{\text{img}_{i}}\}_{i=0}^{n}\in\mathbb{R}^{(h\times w)\times d}$ denote the condition features and the final output of the $i$-th image, respectively. Moreover, the final generated image is denoted as $o_{\text{img}_{n}}\in\mathbb{R}^{(h\times w)\times d}$. As illustrated in the Figure 3 (e), the attention mask in the AR Transformer of MRAR is implemented as a causal attention matrix. The inference process of MRAR is shown in the Appendix.